how many years old is a person if her age now is equal to her age in 3years x 3 less 3 times her age 3 years ago?

Answer by macston(5194)

a=her age
a=3(a+3yr)-3(a-3yr)
a=3a+9yr-3a+9yr
a=18 yrs ANSWER: She is 18 years old
CHECK:
a=3(a+3yr)-3(a-3yr)
18yrs=3(18yrs+3yr)-3(18yrs-3yr)
18yrs=3(21yrs)-3(15yrs)
18yrs=63yrs-45yrs
18yrs=18yrs